1. A woman weighs 129lbf on the earth. a. What does she weigh on the planet Mars? You will need to look up gmas b. c. d. e. f. What is her mass in lbm on both earth and mars? What is her mass in slugs on both earth and mars? What is her mass in grams on both earth and mars? What is her weight in stones (look up stone as an ancient unit) on both earth and mars? What is her weight in metric tons on both earth and mars? 2. A 2-dimensional vector has a length of 25m and makes an angle of 30 deg with the x-axis. a. b. c. d. What is the magnitude of the x-component of the vector? What is the magnitude of the y-component of the vector? What is the angle between the vector and the x-axis in radians? What is the angle between the vector and the y-axis in degrees? 3. A You have three vectors, R- 4 +79+ 62, -9+29 + 112, and š-5R-39-82 a. What is the resultant of these three vectors? b. What is the dot product of R and Q? c. What is the cross product of R and Q? d. What is the mixed triple product of R, Q and S? e. What is the vector Q- S? f. What is the vector R+Q? 4. A triangle has sides of lengths A-6m, B-11m, and C-9m. a. b. c. What is the angle between sides A and B in degrees? What is the angle between sides A and C in degrees? What is the angle between sides B and C in degrees? 5. A gas has a molecular weight of 40g/g-mol. a. Find the specific gas constant R in units of s b. If the absolute pressure is 100lbf/in and the temperature is 150°F, what is the mass ft lbf density of this gas in units of slug/ft?
Answer #1

1. Weight of women on earth = 129 lbf

a) g​​​​​​gmars= 3.8m/s​​​​​2​​​​​s2 = 12.467ft/s​​​​​​2

g​​​​​​earth= 9.8m/s​​​​​​2 = 32.2ft/s2

Weight on mars = mass* acceleration due to gravity of mars

= (129/32.2)*12.467 = 49.95 lbf ~ 50lbf

b) mass on earth and mars will be equal.

Mass (in slug)= (weight in lbf)/(acceleration due to gravity)

= 129/32.2 = 4 slug

1 slug = 32.2 lbm

So mass ~ 129lbm

c) Mass in slug on both earth and mars =4 slug

d) 1slug = 14593.9 g

Mass equal 4 slug = 58375.6 g

2. Legth of 2-d vector =25 unit , angle with x axis =(p)= 30 deg

X-component of vector= 25 Cos(p) = 25 cos30= 21.65m

Y-componet of vector= 25 sin(30)= 12.5m

Angle between x-axis and vector= 30 deg. = 30*(π÷180) = π/6

Angle between y-axis and vector = 60 deg = 60*(π/180) = π/3